# Logit Scaling for Out-of-Distribution Detection

### v1

@@ Abstract
The safe deployment of machine learning and AI models in open-world settings hinges critically on the ability to detect out-of-distribution (OOD) data accurately, data samples that contrast vastly from what the model was trained with. Current approaches to OOD detection often require further training the model, and/or statistics about the training data which may no longer be accessible. Additionally, many existing OOD detection methods struggle to maintain performance when transferred across different architectures. Our research tackles these issues by proposing a simple, post-hoc method that does not require access to the training data distribution, keeps a trained network intact, and holds strong performance across a variety of architectures. Our method, Logit Scaling (LTS), as the name suggests, simply scales the logits in a manner that effectively distinguishes between in-distribution (ID) and OOD samples. We tested our method on benchmarks across various scales, including CIFAR-10, CIFAR-100, ImageNet and OpenOOD. The experiments cover 3 ID and 14 OOD datasets, as well as 9 model architectures. Overall, we demonstrate state-of-the-art performance, robustness and adaptability across different architectures, paving the way towards a universally applicable solution for advanced OOD detection.
@@
out-of-distribution detection, distribution shift, robustness
@@ 1 Introduction
Figure 1: Overview of the LTS method. LTS works at inference time during forward pass. It takes features representations and computes sample-specific scalar value which is then used to scale the logits. Final OOD detection score is calculated by applying scoring function to scaled logits. LTS incurs minimal computational costs and it doesn’t modify activations in any way thus completely preserves original working of the network while enhancing OOD detection significantly.
In the forthcoming era of artificial intelligence (AI), ensuring the safety and reliability of AI models is of utmost importance. One of the key components in achieving this is the effective detection of out-of-distribution (OOD) data on which the trained models usually severely underperform [1, 2]. Such underperformance poses a significant challenge to model reliability and safety. This is particularly crucial in areas such as healthcare, autonomous driving, and finance where the ability to detect OOD instances can profoundly influence outcomes. Thus, ensuring models can accurately recognize and manage OOD samples not only enhances their performance but also plays a vital role in preventing potential errors in critical applications.
The flip side of OOD is in-distribution (ID) data, which usually correspond to the exact training data of the model, or data samples from the same distribution as the training data. AI and machine learning research preceding the current phase of large models rely on standardized training sets that are made open and accessible, but the situation has changed starting from the large models era, where the majority of state-of-the-art models have their training data undisclosed or difficult to access. In light of this change, we are in greater need of OOD detection methods that do not necessarily require full access to the model’s training data.
The line of work in OOD detection has advanced drastically in recent years. ReAct [3] first revealed that the activation patterns in the penultimate layer of networks conceal information valuable for identifying out-of-distribution samples. ReAct [3] achieved strong OOD performance by clipping activations of the penultimate layer utilizing training data statistics. Furthermore, the methods ASH [4] and SCALE [5] demonstrated that adjusting activations in the penultimate layer of the network by pruning and scaling on a per-sample basis leads to strong OOD detection performance without the access to in-distribution data. Both methods delivered robust OOD performance, yet altering activations also resulted in a slight reduction in model accuracy. Moreover, neither of said methods demonstrated successful OOD detection over a diverse set of neural network architectures. Another recent line of work, ATS [6], has demonstrated that applying a per-sample scaling factor, calculated from training data statistics, to scale logits presents an effective method for out-of-distribution detection when applied as an addition to some of the well estabilished methods such as ReAct, ASH, Dice etc.
@@
In this paper, we present a logit scaling (LTS [^1] Code is available at: https://github.com/andrijazz/lts) method for out-of-distribution detection. Our method operates completely post-hoc, meaning that model does not rely on the statistics derived from the training data. This method is ready to use off-the-shelf, offering a remarkably simple, flexible and cost-effective solution. LTS leverages insights from Sun et al. [3], Djurisic et al. [4], Xu et al. [5] and relies on the feature representation from the penultimate layer. Based on that respresentation, it computes a scaling factor on a per-sample basis by using the relationship between strong and week activations. Similarly to ATS [6] the scaling factor is used to scale the logits. Once scaled, logits are then used to compute the final OOD score using well established OOD scoring function, energy score [9]. Figure 1 summarizes the inner workings of our method.
@@
  • We propose a simple and post-hoc method LTS for OOD detection that can be used off-the-shelf without access to in-distribution data.
@@ 2 Related work
The field of out-of-distribution detection has significantly expanded in recent years, driven by the necessity for safer ML model deployment. Baseline methods such as those proposed by Hendrycks and Gimpel [2], Liu et al. [9], and Lee et al. [10] have laid the groundwork for much of the subsequent research in this area. We discuss four major groups of OOD detection methods relevant for understanding our own work.
OOD detection methods leveraging training data statistics. A number of earlier techniques, including ReAct [3], DICE [7], and KNN [11] employ statistics derived from training data in their OOD detection process. Specifically, ReAct adjusts the penultimate layer activations using a clipping threshold calculated from the training data. However, these methods often face challenges such as the unavailability of training data and the necessity for additional post-training processing to extract these statistics. In contrast, the proposed LTS method operates independently of training data statistics and is entirely post-hoc.
Post-hoc per sample treatment for OOD detection. In recent years, as models have become increasingly larger and their training datasets, too, the requirement to access the training data has become impractical, paving the way for methods that do not rely on the training data. ASH [4] found that treatment of activations on per-sample basis enhances OOD detection performance. Instead of implementing a global enhancement based on in-distribution (ID) data, ASH offers specific adjustments tailored to each individual sample. Follow-up work SCALE [5] also adopted this strategy, focusing on the per-sample treatment of penultimate layer activations and improved the activation treatment process. LTS employs a similar approach.
Temperature scaling. Temperature scaling, as introduced by Guo et al. [12], adjusts the logits of a neural network using a scalar value learned from a distinct validation dataset. More recently, Krumpl et al. [6] introduced the ATS, which applies temperature scaling on a per-sample basis. ATS leverages training data statistics, utilizing the activations of each sample to determine an individual scaling factor. While LTS adopts a similar approach, there are two key distinctions: firstly, unlike ATS, our method does not rely on training data statistics; secondly, LTS is an independent detection method, while ATS is used to enhance existing methods such as ReAct, ASH, and DICE.
@@ 3 Logit Scaling for OOD Detection
Our method is inspired by findings reported in ReAct [3], which highlight that out-of-distribution samples often induce abnormally high activations. These findings are illustrated by Figure 2. Several methods tried to exploit this discrepancy of patterns observed in out-of-distribution and in-distribution samples [3, 4, 7, 5]. Our subsequent research, notably ASH-S [4], first observed that treating activations on a sample-based approach enhances the efficacy of OOD detectors and also lifts the requirement of accessing the training set at detection time. Still, it relied on activation pruning which resulted in modest losses of accuracy. Now we make a further step forward by proposing a method which avoids any network modification and relies only on the modification of the popular energy score.
The energy score, as introduced by Liu et al. [9], is one of the most frequently employed techniques for detecting out-of-distribution samples. The process of detecting OOD samples using energy score operates in the following manner: raw inputs are processed by the network, the logits are computed, and fed into the energy score function. This function then produces a score used to classify whether a sample is in-distribution (ID) or out-of-distribution (OOD). The energy score is defined as a scalar value computed by the formula:
@@
3:$f\leftarrow\text{logits}(h)$
4:$S1\leftarrow\sum_{i=1}^{n}h_{i}$
5:$h_{top}=\text{top}_{k}(h,p\cdot n)$
6:$S2\leftarrow\sum_{t\in h_{top}}t$
7:$S\leftarrow\left(\frac{S1}{S2}\right)^{2}$
8:return $S$
Such simple treatment increases the scaling factor $S$ for ID samples and reduces it for OOD samples, effectively influencing separation of two distributions. The sample-based scaling factor $S$ is then used to scale the logits, facilitating the use of the energy score for the final OOD detection as shown in  Figure 1. Figure 3 illustrates the effect of applying LTS on the logits distribution and OOD detection score.
Figure 3: Effect of LTS Treatment. Plots demonstrate the changes in the distribution of logits and Energy scores resulting from LTS treatment. The left-hand plots represent the state before LTS application, while the right-hand plots represent the state after LTS is applied. The plots were generated using a ResNet-50 architecture pretrained on the ImageNet-1k (ID) dataset, with iNaturalist serving as the OOD dataset. The application of LTS produces more extreme logit distribution for OOD samples and improves the separation between ID and OOD scores, leading to a substantial enhancement in OOD detection performance. The logit distribution plots were generated using a single ID and a single OOD sample, whereas the energy score plots were created using 200 images sampled from both the ID and OOD datasets.
@@ 4 Experiments
In this section, we introduce benchmarks, outline evaluation metrics, and present the experimental results of our approach, with a focus on evaluating the effectiveness of LTS and its applicability across different architectures.
@@ 4.3 OOD detection performance
Table 4: LTS performance on OpenOOD benchmark. On the NearOOD task, LTS performs comparably to the previous state-of-the-art and slightly improves it in terms of the FPR@95 evaluation metric. On the FarOOD task, LTS performs similarly to the existing state-of-the-art method. $\uparrow$ indicates larger values are better and $\downarrow$ indicates smaller values are better. All values are percentages. Method results except for LTS (marked as “Ours”) are taken from Xu et al. [5].
@@
4.4 LTS applicability accross architectures
A recent study by Zhao et al. [8] revealed that many previous methods fail to maintain robust out-of-distribution performance across different architectures. Specifically, modern transformer-based architectures such as ViT [30], SwinTransformer [31], as well as MLP-Mixer [32] present significant challenges for existing OOD detection methods. The OptFS [8] method was the first to demonstrate consistent performance across eight different architectures. In our experiments, we replicate the exact experimental setup of OptFS and demonstrate that our method performs consistently well across all tested architectures, significantly reducing FPR@95 while maintaining a comparable level of AUROC compared to OptFS. Figure 4 illustrates the performance of LTS across five different architectures. Detailed results for all tested architectures can be found in Appendix B. Integration of LTS within the different architectures is described in Appendix C.
(a) AUROC evaluation (higher means better)
(b) FPR@95 evaluation (lower means better)
Figure 4: Performance comparison of OOD methods across five architectures. This figure compares the performance of various OOD detection methods across five different architectures. The evaluation is based on two metrics: (a) AUROC and (b) FPR@95. All results are tested on ImageNet benchmark and averaged across 4 tasks (iNaturalist, SUN, Places, and Textures). Higher AUROC indicates better performance, while lower FPR@95 indicates better performance.
@@ 4.5 Ablation studies
Figure 5: Estimating optimal threshold $p$ for scaling factor calculation. We evaluated LTS with various values of the hyperparameter $p$ across five architectures and averaged the results over four tasks. Our findings indicate that LTS achieves optimal performance at $p=5\%$ on both evaluation metrics, AUROC and FPR@95. Therefore, we recommend using $p=5\%$ as the default value.
Determining the optimal value of $p$. In this section we evaluate the performance of LTS across various values of $p$, which represents the proportion of top activations used in computing the scaling factor $S$. The detailed analysis is shown in Figure 5. The appropriate choice of $p$ is highly dependent upon the specific task as well as network architecture in use. In Figure 5, we present a comprehensive sweep across five architectures: ResNet-50, MobileNet-V2, ViT-16/B, Swin Transformer S and MLP-Mixer-B and 4 different tasks from ImageNet-1k benchmark. Based on our analysis, we have found that $p$ value of 5% generally yields optimal results.
Compatibility with other OOD detection methods. LTS is not compatible with methods that are relying on scaling activations in middle layers of the network due to its operational characteristics. For instance, ASH-S and SCALE perform scaling of activations on the penultimate layer, which is quite similar to the LTS operation. Consequently, combining these methods with LTS is not practical.
On the other hand, ReAct [3] greatly benefits from LTS. ReAct leverages the observation illustrated by Figure 2 to clip penultimate layer activations, thus distinguishing between in-distribution and out-of-distribution data. In our approach, we first use LTS to calculate the scaling factor $S$, then apply the ReAct rectification operation, and finally scale the logits using $S$. Table 6 presents the results of combining these two methods, showing that LTS enhances ReAct’s performance across all metrics.
@@
FPR95 AUROC AUPR
$\downarrow$ $\uparrow$ $\uparrow$
@@
Table 6: Compatibility of LTS and ReAct. ReAct significantly benefits when combined with LTS. Our approach involves first using LTS to calculate the scaling factor $S$, then applying the ReAct rectification operation, and finally scaling the logits using $S$. In the results table, $\uparrow$ indicates that higher values are better, while $\downarrow$ signifies that lower values are preferable. Presented results are averaged across 4 tasks on ImageNet benchmark.

### v2

@@ Abstract
The safe deployment of machine learning and AI models in open-world settings hinges critically on the ability to detect out-of-distribution (OOD) data accurately, data samples that contrast vastly from what the model was trained with. Current approaches to OOD detection often require further training the model, and/or statistics about the training data which may no longer be accessible. Additionally, many existing OOD detection methods struggle to maintain performance when transferred across different architectures. Our research tackles these issues by proposing a simple, post-hoc method that does not require access to the training data distribution, keeps a trained network intact, and holds strong performance across a variety of architectures. Our method, Logit Scaling (LTS), as the name suggests, simply scales the logits in a manner that effectively distinguishes between in-distribution (ID) and OOD samples. We tested our method on benchmarks across various scales, including CIFAR-10, CIFAR-100, ImageNet and OpenOOD. The experiments cover 3 ID and 14 OOD datasets, as well as 9 model architectures. Overall, we demonstrate state-of-the-art performance, robustness and adaptability across different architectures, paving the way towards a universally applicable solution for advanced OOD detection. Our code is available at http://github.com/andrijazz/lts.
@@
out-of-distribution detection, distribution shift, robustness, post hoc, generalization
@@ 1 Introduction
Figure 1: Overview of the LTS method for OOD detection. LTS works at inference time during forward pass. It takes features representations and computes sample-specific scalar value which is then used to scale the logits. Final OOD detection score is calculated by applying scoring function to scaled logits. LTS incurs minimal computational costs and it doesn’t modify activations in any way thus completely preserves original working of the network while enhancing OOD detection significantly.
In the forthcoming era of artificial intelligence (AI), ensuring the safety and reliability of AI models is of utmost importance [1, 2, 3]. One of the key components in achieving this is the effective detection of out-of-distribution (OOD) data on which the trained models usually severely underperform [4, 5]. Such underperformance poses a significant challenge to model reliability and safety. This is particularly crucial in areas such as healthcare, autonomous driving, and finance where the ability to detect OOD instances can profoundly influence outcomes. Thus, ensuring models can accurately recognize and manage OOD samples not only enhances their performance but also plays a vital role in preventing potential errors in critical applications.
The flip side of OOD is the in-distribution (ID) data, which usually correspond to the exact training data of the model, or data samples from the same distribution as the training data. AI and machine learning research preceding the current phase of large models rely on standardized training sets that are made open and accessible, but the situation has changed starting from the large models era, where the majority of state-of-the-art models have their training data undisclosed or difficult to access. In light of this change, we are in greater need of OOD detection methods that do not necessarily require full access to the model’s training data.
The line of work in OOD detection has advanced drastically in recent years. ReAct [6] first revealed that the activation patterns in the penultimate layer of the networks conceal valuable information for identifying out-of-distribution samples. ReAct [6] achieved strong OOD performance by clipping activations of the penultimate layer utilizing training data statistics. Furthermore, the methods ASH [7] and SCALE [8] demonstrated that adjusting activations in the penultimate layer of the network by pruning and scaling on a per-sample basis leads to strong OOD detection performance without access to in-distribution data. Both methods delivered robust OOD performance, yet altering activations also resulted in a slight reduction in model accuracy. Moreover, neither of said methods demonstrated successful OOD detection over a diverse set of neural network architectures.
@@
In this paper, we present a logit scaling (LTS) method for out-of-distribution detection. Our method operates completely post-hoc, meaning that model does not rely on the statistics derived from the training data. This method is ready to use off-the-shelf, offering a remarkably simple, flexible and cost-effective solution. LTS leverages insights from Sun et al. [6], Djurisic et al. [7], Xu et al. [8] and relies on the feature representation from the penultimate layer. Based on that respresentation, it computes a scaling factor on a per-sample basis by using the relationship between strong and week activations. The scaling factor is used to scale the logits. Once scaled, logits are then used to compute the final OOD score using well established OOD scoring function, energy score [12]. Figure 1 summarizes the inner workings of our method.
@@
  • We propose a simple and post-hoc method LTS, that can be used off-the-shelf without access to in-distribution data.
@@ 2 Related work
The field of out-of-distribution detection has expanded significantly in recent years, driven by the need for safer ML model deployment. Baseline methods such as those proposed by Hendrycks and Gimpel [5], Liu et al. [12], and Lee et al. [13] have laid the foundation for much of the subsequent research in this area. We discuss four major groups of OOD detection methods relevant for understanding our own work.
Methods relying on training data statistics. Techniques such as ReAct [6], DICE [9], KNN-based methods [14] and ATS [11] use in-distribution data to derive thresholds or calibration parameters. These methods are effective but require access to training data and often involve additional processing steps after training.
Post-hoc model enhancement methods. To address the impracticality of accessing large-scale training data, methods such as ASH [7] and SCALE [8] modify penultimate-layer activations per sample without statistical information derived from training data. Although these approaches improve detection, they can negatively affect the network’s performance on the original task. LTS similarly operates post hoc but avoids modifying activations directly, thereby completely preserving ID performance and leaving network inner working intact. Additional details on the differences between our method and SCALE are provided in Appendix D.
Logit scaling and temperature-based methods. Another family of post-processing approaches works by calibrating the output logits or softmax scores to better distinguish OOD uncertainty. Techniques such as ODIN [15] apply a fixed high-temperature scaling to soften the softmax distribution (often combined with input perturbation), while more recent methods like Adaptive Temperature Scaling (ATS) [11] choose a temperature per sample based on intermediate-layer signals and in combination with empirical cumulative distribution function derived from training data statistics. Further details on the differences between our approach and ATS are provided in Appendix E.
@@ 3 Logit Scaling for OOD Detection
The energy score, as introduced by Liu et al. [12], is one of the most frequently employed techniques for detecting out-of-distribution samples. The process of detecting OOD samples using energy score operates in the following manner: raw inputs are processed by the network, the logits are computed, and fed into the energy score function. This function then produces a score used to classify whether a sample is in-distribution (ID) or out-of-distribution (OOD). In this work, we propose a modification to this standard procedure.
The energy score is defined as a scalar value computed by the formula:
@@
Such simple treatment increases the scaling factor $S$ for ID samples and reduces it for OOD samples, effectively influencing separation of two distributions. The sample-based scaling factor $S$ is then used to scale the logits, facilitating the use of the energy score for the final OOD detection as shown in  Figure 1. Figure 2 illustrates the impact of applying LTS on the logits distribution and the resulting OOD detection scores. A more detailed illustration of the effects of LTS as the hyperparameter $p$ is varied is provided in Appendix F.
Figure 2: Effect of LTS Treatment. Plots demonstrate the changes in the distribution of logits and Energy scores resulting from LTS treatment. The left-hand plots represent the state before LTS application, while the right-hand plots (zoomed-in) represent the state after LTS is applied. The plots were generated using a ResNet-50 architecture pretrained on the ImageNet-1k (ID) dataset, with iNaturalist serving as the OOD dataset. The application of LTS produces more extreme logit distribution for OOD samples and improves the separation between ID and OOD scores, leading to a substantial enhancement in OOD detection performance. The logit distribution plots were generated using a single ID and a single OOD sample, whereas the energy score plots were created using 200 images sampled from both the ID and OOD datasets.
@@
3:$S1\leftarrow\sum_{i=1}^{n}h_{i}$
4:$h_{top}=\text{top}_{k}(h,p\cdot n)$
5:$S2\leftarrow\sum_{t\in h_{top}}t$
6:$S\leftarrow\left(\frac{S1}{S2}\right)^{2}$
7:return $S$
Our method is inspired by findings reported in ReAct [6], which highlight that out-of-distribution samples often induce abnormally high activations. These findings are illustrated by Figure 3. Moreover, in our prior work ASH [7] we observed that high-magnitude activations tend to capture the information most useful for the network’s task. Namely, we showed that it is possible to prune up to 90% of the lowest activations while still maintaining strong model performance. These two observations justify our treatment of high-magnitude activations as particularly important for the out-of-distribution detection task. Therefore, in the LTS we define the scaling factor so as to emphasize the contribution of the strongest activations (the top $p$ percent of $h$) that are capturing the differences in activation patterns between ID and OOD data. Leveraging these discrepancies, we use the scaling factor to adjust the logits, ultimately enhancing the separation between in-distribution and out-of-distribution samples.
@@ 4 Experiments
In this section, we describe the experimental setup used to evaluate our method. We first introduce the benchmarks, covering a diverse range of in-distribution and out-of-distribution datasets across different scales and domains. We then outline the evaluation metrics and finally, we present experimental results that demonstrate the strong performance and generalizability of LTS, with particular attention to its applicability across variety of architectures.
@@ 4.3 OOD detection performance
Table 4: LTS performance on OpenOOD benchmark. The model used for these results is ResNet-50 pretrained on ImageNet-1k (ID) dataset. On the NearOOD task, LTS performs comparably to the previous state-of-the-art and slightly improves it in terms of the FPR@95 evaluation metric. On the FarOOD task, LTS is the second best model in terms of both metrics. $\uparrow$ indicates larger values are better and $\downarrow$ indicates smaller values are better. All values are percentages. Method results except for LTS (marked as “Ours”) are taken from Xu et al. [8].
@@
4.4 LTS applicability across architectures
A recent study by Zhao et al. [10] revealed that many previous methods fail to maintain robust out-of-distribution performance across different architectures. Specifically, modern transformer-based architectures such as Vision Transformer [33], Swin Transformer [34], as well as MLP-Mixer [35] present significant challenges for existing OOD detection methods. The OptFS [10] method was the first to demonstrate consistent performance across eight different architectures. In our experiments, we replicate the exact experimental setup of OptFS and demonstrate that our method performs consistently well across all tested architectures, significantly reducing FPR@95 while maintaining a comparable level of AUROC compared to OptFS. Figure 4 illustrates the performance of LTS across five different architectures. Detailed results for all tested architectures can be found in Appendix B. Integration of LTS within the different architectures is described in Appendix C.
Figure 4: Performance comparison of OOD methods across five architectures. This figure compares the performance of various OOD detection methods across five different architectures. The evaluation is based on two metrics: AUROC (top figure) and FPR@95 (bottom figure). All results are tested on ImageNet-1k benchmark and averaged across 4 tasks (iNaturalist, SUN, Places, and Textures). Higher AUROC indicates better performance, while lower FPR@95 indicates better performance.
4.5 Failure cases and limitations
LTS achieves state-of-the-art results overall but slightly underperforms SCALE on the relatively easy Far-OOD benchmark, where both methods achieve comparable performance. We hypothesize that SCALE’s advantage arises from its aggressive exponential scaling, which can be particularly effective on OOD samples that are substantially different from the in-distribution data. Appendix G presents visualizations of LTS failure cases, illustrating challenges such as subtle domain shifts, texture bias, and high semantic overlap between ID and OOD samples.
Non-convolutional architectures such as Vision Transformers, Swin Transformers, and MLP-Mixers present additional challenges for OOD detection. In these models, the penultimate layer features tend to be more complex and less structured compared to convolutional networks, due to their global receptive fields and attention-based mechanisms. This added complexity makes distinguishing between in-distribution and out-of-distribution samples more difficult. As shown in  Figure 4, we observe slight degradation in performance, specifically an increase in the false positive rate, when applying LTS to non-convolutional architectures. We present LTS failure cases on ViT-B and Swin-B architectures in Appendix G. Addressing these challenges and improving OOD detection on such architectures will be an important direction for our future work.
Apart from these empirically observed challenges, there are other potential limitations of our method and other similar methods. Namely, the energy score relies on meaningful information conveyed via logits. In the ideal case, ID samples will have a large positive logit, say $f_{i}(\mathbf{x})$ for one class, and large negative logits for other classes. In extreme, the energy score is close to $-f_{i}(\mathbf{x})$. On the other hand, OOD samples will have all negative logits (since no class is the appropriate one). In extreme, the energy score is a large positive number. However, these properties depend on the quality of the analyzed model, specifically the quality of its penultimate layer activations. For instance, if the model confuses an OOD sample for a representative of some class and assigns it a high logit, the energy score might be low or negative and thus deceiving. If the model cannot separate some classes, their probabilities tend to be conservative (in order to avoid large errors which the cross-entropy penalizes severely), resulting in low logits, which might lead to higher values of the energy score even for ID samples. In conclusion, various practical obstacles during training, such as noisy data or extreme class imbalance, can reduce the quality of the analyzed model. These might also negatively impact the OOD detection performance of energy-score-based methods (including ours) for that model.
@@ 4.6 Ablation studies
Determining the optimal value of $p$. We examine how varying the hyperparameter $p$, which specifies the proportion of top activations used to compute the scaling factor $S$, affects the performance of our method.
The optimal value of $p$ depends on both, the underlying architecture and ID dataset. To assess the effect of $p$, we conduct a comprehensive evaluation across five architectures: ResNet-50, MobileNet-V2, ViT-B/16, Swin Transformer-S, and MLP-Mixer-B, using four tasks from the ImageNet-1k benchmark. Additionally, we perform a complementary experiment in which the architecture is fixed to DenseNet, and performance is evaluated across three different in-distribution datasets: CIFAR-10, CIFAR-100, and ImageNet-1k. Across all experiments, our empirical results consistently indicate that setting $p=5\%$ yields strong performance on all metrics. A detailed illustration is presented in Figure 5.
Estimating optimal threshold $p$ across different architectures
Estimating optimal threshold $p$ across different ID datasets on DenseNet architecture
Figure 5: Analysis of optimal $p$ values for scaling factor calculation. The top row shows the performance of LTS evaluated across five architectures using various values of the hyperparameter $p$, with results averaged over four tasks. The bottom row focuses on a single architecture, namely DenseNet, pretrained on three different in-distribution datasets: CIFAR-10, CIFAR-100, and ImageNet-1k. Across all our experiments, LTS consistently achieves optimal performance at $p=5\%$ on both evaluation metrics, AUROC and FPR@95. We therefore recommend using $p=5\%$ as the default setting.
Compatibility with other OOD detection methods. LTS is not compatible with methods that scale or prune activations in intermediate layers, such as ASH-S [7] and SCALE [8], due to its operational characteristics.
In contrast, ReAct [6] complements LTS effectively. ReAct clips penultimate layer activations to distinguish in-distribution from out-of-distribution data. In our approach, we first compute the scaling factor $S$ using LTS, apply ReAct’s clipping, and finally scale the logits using $S$. Table 6 shows that this combination consistently improves ReAct’s performance across all evaluation metrics.
@@
FPR95 $\downarrow$ AUROC $\uparrow$ AUPR $\uparrow$
@@
Table 6: Compatibility of LTS and ReAct. ReAct significantly benefits when combined with LTS. Our approach involves first using LTS to calculate the scaling factor $S$, then applying the ReAct rectification operation, and finally scaling the logits using $S$. Presented results are averaged across 4 tasks on ImageNet benchmark.